TO HW 4 Help Save&Exlt Submit Check my work Marvel Parts, Inc., manufactures auto accessories. One of the companys products is a set of seat covers that can be adjusted to fit nearly any small car. The company has a standard cost system in use for all of its products. According to the standards chat have been set for the seat covers, the factory should work 2.850 hours each month to produce 1,900 sets of covers. The standard costs associated with this level of production are 10 points Set Total/ll/of Covers eBook Print References Direct materials Direct labor Variable manufacturing overhead (based 42,560$22-40 S1,30027.00 on direct labor-houre) S 6.840N.60 53.00 During August, the factory worked only 2.800 direct labor-hours and produced 2,000 sets of covers. The following actual costs were recorded during the month: Per Set ocaof Covera 4S,00 $22.80 Dlroct naterisie 13,000 asds23 2 Direct Labor Variable manufacturing overhoad 49.00024.50 s.000W3.5 50.80 At standard, each set of covers should require 5.6 yards of material. All of the materials purchesed during the month were usedin Next < Prev 4 of4
media%2Fe87%2Fe872da54-15da-4006-b420-c2
Answer #1
1 Material price variance=Actual material used*(Standard rate for material-Actual rate for material)
Actual material used=12000 yards
Standard rate=Cost per set of covers/Material required per set of covers=22.40/5.6=$4 per yard
Actual rate=45600/12000=$3.8 per yard
Material price variance=12000*(4-3.8)=2400 F
(Actual rate is less than standard rate.Hence variance is favorable)
Material quantity variance=Standard rate for materials*(Standard materials required-Actual material used)
Standard rate=Cost per set of covers/Material required per set of covers=22.40/5.6=$4 per yard
Standard materials required=Actual set of covers produced*Material required per set of covers=2000*5.6=11200 yards
Actual material used=12000 yards
Material quantity variance=4*(11200-12000)=3200 U
(Actual material used is more than the standard material required.Hence, variance is unfavorable)
2 Labor rate variance=Actual hours worked*(Statndard labor rate-Actual labor rate)
Actual hours worked=2800 hours
Standard labor rate=Standard direct labor cost/Budgeted hours=51300/2850=$18 per hour
Actual labor rate=Actual direct labor cost/Actual hours=49000/2800=$17.5 per hour
Labor rate variance=2800*(18-17.5)=1400 F
(Actual labor rate is less than standard labor rate.Hence variance is favorable)
Labor efficiency variance=Standard labor rate*(Standard labor hours required-Actual labor hours worked)
Standard labor rate=Standard direct labor cost/Budgeted hours=51300/2850=$18 per hour
Standard labor hours required=Actual set of covers produced*Labor hours required per set of covers
Labor hours required per set of covers=Budgeted labor hours/Budgeted production of set of covers=2850/1900=1.5 hours per set of covers
Standard labor hours required=2000*1.5=3000 hours
Actual hours worked=2800 hours
Labor efficiency variance=18*(3000-2800)=3600 F
(Actual hours worked is less than the standard hours required.Hence, variance is favorable)
3 Variable overhead rate variance=Actual hours worked*(Statndard variable overhead rate-Actual variable overhead rate)
Actual hours worked=2800 hours
Standard variable overhead rate=Standard variable overhead cost/Budgeted hours=6840/2850=$2.4 per hour
Actual variable overhead rate=Actual variable overhead cost/Actual hours=7000/2800=$2.5 per hour
Variable overhead rate variance=2800*(2.4-2.5)=280 U
(Actual variable overhead rate is more than standard variable overhead rate.Hence variance is unfavorable)
Variable overhead efficiency variance=Standard variable overhead rate*(Standard labor hours required-Actual labor hours worked)
Standard variable overhead rate=Standard variable overhead cost/Budgeted hours=6840/2850=$2.4 per hour
Standard labor hours required=3000 hours (as computed in part 2)
Actual hours worked=2800 hours
Variable overhead efficiency variance=2.4*(3000-2800)=480 F
(Actual hours worked is less than the standard hours required.Hence, variance is favorable)